OpenStudy (anonymous):

Simplify the Expression cot^2 x+sin^2 x/co^t2 x−csc^2 x Answer Choices: csc x -1 1 sec x

OpenStudy (anonymous):

cot^2 x+sin^2 x/cot^2 x−csc^2 x

OpenStudy (anonymous):

dixed the question

OpenStudy (anonymous):

*fixed

OpenStudy (anonymous):

let me give you the solution in microsoft word

OpenStudy (anonymous):

but i kind of wanted explained as well like how i mean if you can

OpenStudy (anonymous):

first do some algebra that has nothing to do with trig if you call \(\cos(x)=a\) and \(\sin(x)=b\) then you have \[\frac{\frac{a^2}{b^2}+b^2}{\frac{a^2}{b^2}-\frac{1}{b^2}}\]

OpenStudy (anonymous):

okay so sin^2 x/csc^2 x

OpenStudy (anonymous):

clear the compound fraction by multiplying top and bottom by \(b^2\)

OpenStudy (anonymous):

Sorry which would be the compound fraction? a^2/b^2?

OpenStudy (anonymous):

oh wait um a^2+b^4/a^2-1?

OpenStudy (anonymous):

before we continue, i am going to guess that there is still a typo in your question, because it is none of your possible answers

OpenStudy (anonymous):

yes, you would get \[\frac{a^2+b^4}{a^2-1}\]

OpenStudy (anonymous):

hmm ill write it out exactly

OpenStudy (jdoe0001):

\(\bf \cfrac{cot^2 (x)+sin^2(x)}{cot^2(x)−csc^2(x)} \ \ \ ?\)

OpenStudy (anonymous):

\[\frac{ \cot ^2x+\sin ^2x \frac }{ \cot ^2x-\csc ^2x }\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

jdoe0001 thats it

OpenStudy (anonymous):

\[\frac{\cot^2 (x)+\sin^2(x)}{\cot^2(x)−\csc^2(x)} \]

OpenStudy (anonymous):

Would i use the common identities?

OpenStudy (anonymous):

okay then use the identities?

OpenStudy (anonymous):

that was wrong, should be \[\frac{\cos^2(x)+\sin^4(x)}{\cos^2(x)-1}\]

OpenStudy (anonymous):

i still don't get one of your answer, maybe @jdoe0001 has it

OpenStudy (jdoe0001):

hmm heheh, not much either :(

OpenStudy (jdoe0001):

can you post a screenshot of the material?

OpenStudy (anonymous):

you can write \(\cos^2(x)-1=-\sin^2(x)\) but that doesn't get your answer i have seen a problem almost identical to this where the answer was \(-1\) but this one is not

OpenStudy (anonymous):

But what if I were to use some Identities? Or is that what you've been trying?

OpenStudy (anonymous):

like cos x = 1/sec x

OpenStudy (anonymous):

algebra first identities second it is not one of your possible answers

OpenStudy (anonymous):

@jdoe0001 ill try and post a pic but of the question or the lesson?

OpenStudy (anonymous):

question will suffice

OpenStudy (anonymous):

okay one sec would i need a photo website?

OpenStudy (anonymous):

post a screen shot should work

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

does that work?

OpenStudy (anonymous):

\[\frac{\cos^2 (x)+\sin^2(x)}{\cot^2(x)−\csc^2(x)}\]

OpenStudy (anonymous):

you had \(\cot^2(x)\) up top

OpenStudy (anonymous):

oh haha wow im so sorry that is my bad i wasted so much of your time.

OpenStudy (anonymous):

ill stick to pics from now on

OpenStudy (anonymous):

well lets cut to the chase numerator is \(1\) by the most fundamentalist of all identities \(\cos^2(x)+\sin^2(x)=1\)

OpenStudy (anonymous):

denominator is \(-1\) because if you take \[\cos^2(x)+\sin^2(x)=1\] and divide all by \(\sin^2(x)\) you get \[\cot^2(x)+1=\csc^2(x)\]

OpenStudy (anonymous):

making \(\cot^2(x)-\csc^2(x)=-1\) final answer is \(\frac{1}{-1}=-1\)

OpenStudy (anonymous):

thank you so much for your time satellite have a medal